1.     What would the expenditure multiplier be in an econ[1]omy without government spending or taxes where the MPC is 0.8 and the MPm is 0? Where the MPm is 0.1? Where the MPm is 0.9? Explain why the multiplier might even be less than 1. 



1
Expert's answer
2021-12-09T09:33:20-0500

Expenditure multiplier"=\frac{1}{[1-(MPC-MPM)]}"

"MPC=0.8"

"MPM=0"

Multiplier"=\frac{1}{1-(0.8-0)}=5"

"MPM=0.1:"

Multiplier"=\frac{1}{1-(0.8-0.1)}=3.33"

"MPM=0.9"

Multiplier"=\frac{1}{1-(0.8-0.9)}=0.91"

The multiplier might even be less than 1 because government spending increases are supposed to be financed and hence this comes up with a negative wealth effect.